OpenStudy (anonymous):

if sin θ = 12/13 and cos θ < 0, what is tan θ?

Nnesha (nnesha):

\[\rm sin = \frac{ opposite }{ hypotenuse }~~~~ \cos = \frac{ adjacent }{ hypotenuse } ~~\tan= \frac{ opposite }{ adjacent }\]

Nnesha (nnesha):

make a right triangle  sin is opposite over hypotenuse 12/13=opposite /hypotenuse and then use Pythagorean theorem to find 3rd side of right triangle \[\huge\rm a^2 +b^2 =c^2\]

OpenStudy (♪chibiterasu):

We can set up a triangle and find the missing side. Recall SOH-CAH-TOA. SOH = sin = opposite/hypotenuse We can use the pythagorean theorem to find the value of the other side.  \[a^2 + b^2 = c^2\] \[12^3 + 13^2 = c^2\] \[\sqrt{c^2}\]